A bag contains 3 purple marbles and 4 white marbles. Two marbles are drawn at random. One marble is drawn and not replaced. Then a second marble is drawn. What is the probability that the first marble is white and the second one is purple?
Answer by stanbon(75887)

A bag contains 3 purple marbles and 4 white marbles. Two marbles are drawn at random. One marble is drawn and not replaced. Then a second marble is drawn. What is the probability that the first marble is white and the second one is purple?
-----------
(4/7)(3/6) = 12/42 = 2/7
============================
Cheers,
Stan H.